# Breaking Eternal Inflation: Empirical Viability of a Spontaneous Collapse Scenario

### v1

@@ 1 Introduction
According to quantum field theory, the causality conditions constrain the form of the commutation relations of the quantum field operators, and, as stated in [10, p. 384], when one splits the metric into background and perturbation parts, one faces the issue of which is the space-time defining the causal structure to be used in that construction. On the one hand, one could simply rely on the causal structure of the background metric, which is well defined, as that sector is not subject to a quantum treatment. However, this does not seem like a reasonable path, if one really wants to take the space-time structure to be that associated with the “full metric”. On the other hand, one might be inclined to rely on the “full space-time metric,” but then would face the problem that it is a quantum object and thus for general states no clear structure can be assigned. As mentioned before, in usual treatments of early-universe cosmology, it is customary to quantize both the matter and the space-time perturbations without much attention to the issue, and relying in practice on the first alternative. However, if we aim to obtain a conceptually clear explanatory account, it would not be “legitimate” just to ignore this aspect.
@@
It is worthwhile to note that as a byproduct of this approach, one obtains a natural account for a strong suppression of the generation of primordial gravity waves [22] with respect to expectations of the standard treatments of the issue [23], a result which is, of course, in a much better agreement with the empirical constraints on the matter [24].
@@
In the first case, we would not face a problem; however, the second leads to two types of situations 2a) the zero mode of the inflaton moves backward (in the direction of an increasing potential), or 2b) the mode moves forward (along a direction of decreasing potential). The issue is that when the inflaton is displaced in the direction of an increasing potential, the expansion rate increases, causing these regions to grow faster than regions where the inflaton moves in the opposite direction. Thus, over time, the regions with a higher expansion rate will cover a larger portion of the universe, and eventually dominate it as a whole, and therefore inflation would never end. This is the eternal inflation problem in a nutshell, see for instance [25, 26].
@@
As noted in the standard treatments, this problem is, in fact, tied to the erroneous identification of quantum uncertainties with stochastic fluctuations. Once this point is recognized, the problem simply goes away  [^2] It is worth noting that this misconception regarding fluctuations appears in many branches of physics, including high-energy physics. In a recent work [27], it was shown that the seeming discrepancy between the most recent bounds on the magnitude of the neutron electric dipole moment and the direct estimates of the corresponding quantum uncertainties can be traced to an erroneous interpretation of such quantum uncertainties as statistical fluctuations.. However, as we also mentioned, the standard accounts lack a mechanism to break the homogeneity and isotropy of the quantum state, and thus one cannot possibly account for the emergence of cosmic structure. On the other hand, once we incorporate a spontaneous collapse theory into the picture, we face a context that does include stochastic fluctuations (affecting the dynamics of the inflaton field), a fact which now raises the possibility of the eternal inflation problem.
@@
The manuscript is structured as follows: Section 2 begins with a description contrasting the treatments of the primordial perturbations from the standard and the spontaneous collapse approaches. Then, the detailed quantum treatment is presented in the context of the CSL theory together with semiclassical gravity. To connect with the root of the problem, Section 3 is dedicated to the eternal inflation problem as viewed from within the spontaneous collapse framework. Afterward, in Section 4, we present the analysis to assess the observational viability of our model. Finally, Section 5 is devoted to present the results and conclusions.
2 Primordial perturbations: Comparing the standard and the collapse model
The main objective of this section is to present a comparison of the collapse approach derivation of the primordial power spectrum with that of the standard approach. We summarize the key technical aspects of our approach, which is based on the CSL model in the context of semiclassical treatment of the inflationary universe. In particular, we demonstrate how the interplay between semiclassical gravity and the spontaneous collapse mechanism gives rise to the primordial inhomogeneities.
@@ 2.1 Standard treatment
$\langle{0}|\mathcal{R}_{k}(\eta)\mathcal{R}_{k^{\prime}}(\eta)|\rangle{0}=\int d^{3}k\,e^{i\vec{k}\cdot(\vec{x}-\vec{y})}\,\langle\mathcal{R}(x)\mathcal{R}(y)\rangle.$ (20)
@@ 2.2 Semiclassical treatment: The CSL mechanism and the primordial power spectrum
In this section, we discuss the modification of the standard treatment with one that is based on semi-classical gravity and the quantum dynamics of the collapse theory. This leads to the replacement of the classical field $\delta\phi_{\textbf{k}}$ with the quantum expectation value (also classical) $\langle\hat{\delta}\phi_{\textbf{k}}\rangle$  [^7] In principle, when relying on the semi-classical gravity framework, one should simply and directly replace the classical energy momentum tensor with the corresponding expectation value. However, as shown in [17], the fact that the creation and annihilation operators associated with the zero mode, as well as those corresponding to the spatially dependent modes, commute, together with the assumption that the modes are unentangled (to lowest order), indicates that we may proceed to a treatment in which the fields themselves are replaced by their expectation values. Note, by the way, that for the Bunch Davies vacuum we have $\langle\hat{\delta}\phi_{\textbf{k}}\rangle=0$ and it is only due to the collapse dynamics that the state of the quantum field is modified and this quantity acquires a non-vanishing value.. As a result of all this, we have:
@@
The terms $\overline{\langle\hat{y}_{\textbf{k}}^{2}\rangle}$ and $(\textrm{Re}[A_{k}(\eta)])^{-1}$, appearing in the later expression of the power spectrum, can be obtained by solving explicitly the corresponding CSL equations. We have included such a procedure in Appendix A for the interested reader.
@@ 3.1 Observational viability of the CSL model without restrictions
At this point, it is useful to discuss some relevant numerical values for the main physical quantities appearing in the model. We assume that inflation lasts 65 e-folds approximately, so $a_{\rm end}=a_{\rm ini}e^{65}$. Additionally, we also assume that inflation ends at a temperature of $10^{-4}\text{M}_{P}$; comparing this scale with the CMB temperature today $T_{\rm cmb}\simeq 0.23$ meV, we have $a_{0}/a_{\rm end}=T_{\rm end}/T_{\rm cmb}=10^{-27}$. Therefore, using the previous estimates, we obtain an approximate value for the scale factor at the beginning of inflation $a_{\rm ini}/a_{0}=10^{-56}$. Finally, using the form of the scale factor $a(\eta)=-1/(H\eta)$ with $H\simeq 10^{-9}\text{M}_{P}$ approximately constant during inflation and the normalization $a_{0}=1$, we can obtain an estimate for the conformal time at the beginning of inflation $\tau=1/Ha_{\rm ini}$, which yields $\tau\simeq 10^{8}$ Mpc. Regarding $\lambda_{0}$, during the computation of $P(k)$ in Appendix A, we have used the approximation $\lambda_{0}|\tau|\gg 1$. Hence, we fix $\lambda_{0}=10^{-5}$ Mpc-1 (or in international units $\lambda_{0}=10^{-19}$ s-1), in this way $\lambda_{0}|\tau|=10^{3}$.
@@
The oscillations observed in the power spectrum, Fig. 1, are a characteristic feature of the CSL model and are absent in the standard framework. They arise from the way we relate the primordial spectrum to the (squared) quantum expectation value of each field mode, i.e. $P(k)\propto\langle\delta\phi_{k}\rangle^{2}$; note that this is not the quantum uncertainty in the vacuum, $\langle O^{2}\rangle$, which is the quantity usually associated with the power spectrum in the traditional treatments of inflation. From the computational point of view, the oscillations originate from solving the CSL equations with the choice of initial conditions for the usual modes in BD vacuum. [^11] As shown in Appendix A, the initial condition of the CSL model enters through the constants $C_{1}$ and $C_{2}$ in Eq. (53) leading to Eq. (54). In particular, because the Bunch-Davies (BD) vacuum is imposed together with the CSL evolution, the constants $C_{1}$ and $C_{2}$ do not vanish, so the oscillatory terms in (54) persist. This behavior contrasts with standard treatments of inflation, where one would have $C_{1}=C_{2}=0$ [1], and where the oscillatory contributions are absent. Furthermore, still for $\beta=0$, the function $C(k)$ in (38), which encodes the CSL contribution to the power spectrum, can be approximated as,
@@ 5 Summary and Conclusions
Appendix A Solving the CSL equations in the primordial power spectrum

### v2

@@ 1 Introduction
According to quantum field theory, the causality conditions constrain the form of the commutation relations of the quantum field operators, and, as stated in [81, p. 384], when one splits the metric into background and perturbation parts, one faces the issue of which is the space-time defining the causal structure to be used in that construction. On the one hand, one could simply rely on the causal structure of the background metric, which is well defined, as that sector is not subject to a quantum treatment. However, this does not seem like a reasonable path, if one really wants to take the space-time structure to be that associated with the “full metric”. On the other hand, one might be inclined to rely on the “full space-time metric,” but then would face the problem that it is a quantum object and thus for general states no clear structure can be assigned. As mentioned before, in usual treatments of early-universe cosmology, it is customary to quantize both the matter and the space-time perturbations without much attention to the issue, and relying in practice on the first alternative. However, if we aim to obtain a conceptually clear explanatory account, it would not be “legitimate” just to ignore this aspect.
@@
It is worthwhile to note that as a byproduct of this approach, one obtains a natural account for a strong suppression of the generation of primordial gravity waves [46] with respect to the standard predictions of traditional inflationary models [59].
For relatively large class of specific inflationary models the predictions concerning the spectrum of the tensor modes or primordial gravity waves are so close to the predictions concerning the scalar modes that consideration of the observed value for the amplitude and tilt of the latter together with the absence of detection of the former results in the conclusion that those specific models are ruled out [3, 4];particularly the simplest ones. Although several concrete inflationary models remain viable despite these considerations; in particular, plateau-like inflaton potentials, which are considered unnatural by some theorists [41, 40]. On the other hand, our approach, based on semiclassical gravity together with spontaneous reduction, generically leads to a strong suppression of the amplitude of tensor perturbations as compared to that of scalar perturbations. Thus, most natural inflationary potentials (e.g. $V\sim\phi^{2}$) remain, in principle, viable when analyzed within such a framework.
@@
In the first case, we would not face a problem; however, the second leads to two types of situations 2a) the zero mode of the inflaton moves backward (in the direction of an increasing potential), or 2b) the mode moves forward (along a direction of decreasing potential). The issue is that when the inflaton is displaced in the direction of an increasing potential, the expansion rate increases, causing these regions to grow faster than regions where the inflaton moves in the opposite direction. Thus, over time, the regions with a higher expansion rate will cover a larger portion of the universe, and eventually dominate it as a whole, and therefore inflation would never end. This is the eternal inflation problem in a nutshell, see for instance [51, 38]. As noted in the standard treatments, this problem is, in fact, tied to the erroneous identification of quantum uncertainties with stochastic fluctuations.
That issue, and its connection with the quantum measurement problem, has been emphasized by multiple authors (see, e.g., [61, 83]), and several proposals aimed at sharpening the standard account can be found in the literature. In particular, there is an extensive program addressing the emergence of an effectively classical description during inflation via decoherence and related mechanisms [18, 19], as well as analyses based on backreaction and stochastic effects [20, 53], and discussions stressing the role of vacuum squeezing in cosmology (see, e.g., [9]).
Nevertheless, as discussed in [16, 15] (see also [64, 26]), we do not consider these mechanisms to provide, by themselves, a fully satisfactory account of the specific issue emphasized here [^2] The issue is not in the calculations of the reduced density matrices that underlay decoherence or the analysis of the “squeezzing” of the state of quantum fields tied to the rapid cosmological expansion, but to the fact that these aspects by themselves can not justify the interpretation of the situation as involving any breakdown of the relevant symmetry. In fact as long as we take quantum theory to be more fundamental than the classical one, and thus accepting the classical description superveens on the quantum one, its simply a logical mistake to try to account for a breakdown of symmetry according to the former while admitting it is preserved according to the latter, see for a detailed discussion[15].: the actual breaking of homogeneity and isotropy starting from an exactly homogeneous and isotropic quantum state evolving under symmetric dynamics. While decoherence can render interference terms in reduced density matrices negligible in an appropriate basis, and squeezing might motivate an effective classical description, neither step alone selects a single inhomogeneous realization (i.e. a definite outcome) or justifies interpreting the situation as an actual symmetry-breaking event. In our view, this is precisely where an objective-collapse dynamics becomes relevant. Furthermore, the stochastic element often (and, in our view, incorrectly) attributed to quantum fluctuations is not provided by decoherence or by vacuum squeezing alone.
@@
The latter point is particularly relevant in discussions of eternal inflation, where stochasticity plays a central role, but upon recognition that the traditional approaches provide no genuine source of stochastic behavior (unless some sort of “measurement process” is invoked), the problem would seem to just disappear [^3] It is worth noting that this misconception regarding fluctuations appears in many branches of physics, including high-energy physics. In a recent work [36], it was shown that the seeming discrepancy between the most recent bounds on the magnitude of the neutron electric dipole moment and the direct estimates of the corresponding quantum uncertainties can be traced to an erroneous interpretation of such quantum uncertainties as statistical fluctuations.. However, as we also mentioned, the standard accounts lack a mechanism to break the homogeneity and isotropy of the quantum state, and thus one cannot possibly account for the emergence of cosmic structure. On the other hand, once we incorporate a spontaneous collapse theory into the picture, we face a context that does include stochastic fluctuations (affecting the dynamics of the inflaton field), a fact which now raises the possibility of the eternal inflation problem.
@@
The manuscript is structured as follows: Section 2 begins with a description contrasting the treatments of the primordial perturbations from the standard and the spontaneous collapse approaches. Then, the detailed quantum treatment is presented in the context of the CSL theory together with semiclassical gravity. To connect with the root of the problem, Section 3 is dedicated to the eternal inflation problem as viewed from within the spontaneous collapse framework. Afterward, in Section 4, we present the analysis to assess the observational viability of our model. Finally, Section 5 is devoted to presenting the results and conclusions.
2 Primordial perturbations: comparing the standard and the CSL model
The main objective of this section is to present a comparison of the study of primordial perturbations of the power spectrum from two perspectives: using the standard approach and the Continuous Spontaneous Localization CSL theories. In the first subsection, we will present the standard idea of primordial perturbations, evidencing the conceptual problems at the end. Then, in section 2.2, we presented the study of the primordial power spectrum in light of the CSL theory as has been study in [21, 47]. In particular, we demonstrate how the interplay between semiclassical gravity and the CSL (in the context of semiclassical treatment of the inflationary universe) gives rise to the primordial inhomogeneities. To keep the natural flow of the lecture, we will add the main details of the CSL theory in the appendix A.
@@ 2.1 Standard treatment
$\langle{0}|\mathcal{R}_{k}(\eta)\mathcal{R}_{k^{\prime}}(\eta)|0\rangle=\int d^{3}k\,e^{i\vec{k}\cdot(\vec{x}-\vec{y})}\,\langle\mathcal{R}(x)\mathcal{R}(y)\rangle.$ (20)
@@ 2.2 Semiclassical treatment: The CSL mechanism and the primordial power spectrum
In this section, we summarize the main ideas of modifying the standard treatment based on CSL theory and semi-classical gravity [32, 21, 73]. The details of CSL theory are covered in the Appendix A. CSL theories introduce a change in the dynamics introducing a stochastic term in the evolution equation of the state, this leads to the replacement of the classical field $\delta\phi_{\textbf{k}}$ with the quantum expectation value (also classical) $\langle\hat{\delta}\phi_{\textbf{k}}\rangle$ [^8] In principle, when relying on the semi-classical gravity framework, one should simply and directly replace the classical energy momentum tensor with the corresponding expectation value. However, as shown in [32], the fact that the creation and annihilation operators associated with the zero mode, as well as those corresponding to the spatially dependent modes, commute, together with the assumption that the modes are unentangled (to lowest order), indicates that we may proceed to a treatment in which the fields themselves are replaced by their expectation values. Note, by the way, that for the Bunch Davies vacuum we have $\langle\hat{\delta}\phi_{\textbf{k}}\rangle=0$ and it is only due to the collapse dynamics that the state of the quantum field is modified and this quantity acquires a non-vanishing value.. As a result of all this, we have:
@@
The terms $\overline{\langle\hat{y}_{\textbf{k}}^{2}\rangle}$ and $(\textrm{Re}[A_{k}(\eta)])^{-1}$, appearing in the later expression of the power spectrum, can be obtained by solving explicitly the corresponding CSL equations. We have included such a procedure in Appendix B for the interested reader.
@@ 3.1 Observational viability of the CSL model without restrictions
At this point, it is useful to discuss some relevant numerical values for the main physical quantities appearing in the model. We assume that inflation lasts 65 e-folds approximately, so $a_{\rm end}=a_{\rm ini}e^{65}$. Additionally, we also assume that inflation ends at a temperature of $10^{-4}\text{M}_{P}$; comparing this scale with the CMB temperature today $T_{\rm cmb}\simeq 0.23$ meV, we have $a_{0}/a_{\rm end}=T_{\rm end}/T_{\rm cmb}=10^{-27}$. Therefore, using the previous estimates, we obtain an approximate value for the scale factor at the beginning of inflation $a_{\rm ini}/a_{0}=10^{-56}$. Finally, using the form of the scale factor $a(\eta)=-1/(H\eta)$ with $H\simeq 10^{-9}\text{M}_{P}$ approximately constant during inflation and the normalization $a_{0}=1$, we can obtain an estimate for the conformal time at the beginning of inflation $\tau=1/Ha_{\rm ini}$, which yields $\tau\simeq 10^{8}$ Mpc. Regarding $\lambda_{0}$, during the computation of $P(k)$ in Appendix B, we have used the approximation $\lambda_{0}|\tau|\gg 1$. Hence, we fix $\lambda_{0}=10^{-5}$ Mpc-1 (or in international units $\lambda_{0}=10^{-19}$ s-1), in this way $\lambda_{0}|\tau|=10^{3}$.
@@
The oscillations observed in the power spectrum, Fig. 1, are a characteristic feature of the CSL model and are absent in the standard framework. They arise from the way we relate the primordial spectrum to the (squared) quantum expectation value of each field mode, i.e. $P(k)\propto\langle\delta\phi_{k}\rangle^{2}$; note that this is not the quantum uncertainty in the vacuum, $\langle O^{2}\rangle$, which is the quantity usually associated with the power spectrum in the traditional treatments of inflation. From the computational point of view, the oscillations originate from solving the CSL equations with the choice of initial conditions for the usual modes in BD vacuum. [^12] As shown in B, the initial condition of the CSL model enters through the constants $C_{1}$ and $C_{2}$ in Eq. (55) leading to Eq. (56). In particular, because the Bunch-Davies (BD) vacuum is imposed together with the CSL evolution, the constants $C_{1}$ and $C_{2}$ do not vanish, so the oscillatory terms in (56) persist. This behavior contrasts with standard treatments of inflation, where one would have $C_{1}=C_{2}=0$ [21], and where the oscillatory contributions are absent. Furthermore, still for $\beta=0$, the function $C(k)$ in (38), which encodes the CSL contribution to the power spectrum, can be approximated as,
@@ 5 Summary and Conclusions
Appendix A CSL in a nutshell
Collapse theories have been designed to address the measurement problem in Quantum Mechanics. We remind the reader that this problem lies in the fact that quantum theory, as presented in textbooks, invokes two different evolution rules [69], one completely deterministic, embodied by Schrödinger’s equation, and another one of stochastic nature, that is supposed to apply when a measurement is performed on the quantum system. The issue is that the theory is, in general, quite obscure regarding the general rules determining when one or the other evolution laws apply, that is, what exactly counts as a measurement [^18] Simple proposals such as “all interactions are measurements” are ruled out by interference experiments.. The difficulty becomes apparent when considering Bell’s rhetorical question [12]: “What exactly qualifies some physical systems to play the role of “measurer”? Was the wavefunction of the world waiting to jump for thousands of millions of years until a single-celled living creature appeared?. . .” In fact, this could be most alarming, adding questions like: when does a particular measurement take place?, what quantity is being measured?, etc.
Collapse theories work by unifying the two evolution rules and removing any “apriori” distinction in the manner of treating the system of interest and the apparatus; all the physical systems are treated in quantum mechanical terms. The “special” role of apparatus emerges (rather than being postulated) from the universal evolution law as a result of the large number of degrees of freedom in it, and the entanglement between these and those of the system of interest caused by their interactions. The basic idea is that the effect of the modified general dynamics, involving departure from unitary, differs in a negligible way when a very small number of elementary degrees of freedom are involved, but might become dominant when a vast number of degrees of freedom are involved. This process is called the amplification mechanism, and is a generic feature of the collapse models.
Collapse theories have been originally constructed for applications in the non-relativistic many-particle quantum mechanics. The first realistic version of these theories is the GRW proposal [35], which involves discrete spontaneous reduction of the wave function for the particle’s position degrees of freedom. A very attractive alternative is the Continuous Spontaneous Localization [68], which naturally allows the treatment of identical particles, and involves a Wiener-like continuous reduction process.
It is an important remark that, in the literature, there exist different versions of the same CSL theory; many of them (the non-relativistic cases) are described in [68]. As one might expect, the general idea is the same, but some of the main differences concern the choice of the collapse operator and the collapse parameter. The “most general” version of the theory (in the sense that it allows for a generic form of the collapse operator and its associated collapse rate) is presented in [68].
The mathematical framework consists of two equations: i) a stochastic differential evolution equation, whose solution takes the form:
$|\psi,t\rangle_{w}=\hat{\cal T}e^{-\int_{0}^{t}dt^{\prime}\big[i\hat{H}+\frac{1}{4\lambda}[w(t^{\prime})-2\lambda\hat{A}]^{2}\big]}|\psi,0\rangle.$}} (49)
where $\hat{\cal T}$ s the time-ordering operator, $A$ the collapse operator, $w(t)$ the random classical function of time.
The second equation is the probability rule for the realization $\tilde{w}(t)$ of an element of stochastic class $\{w(t)\}$, restricted to lie in the “tube” in the plane wtw-t defined by the tolerance intervals $\text{d}w(t_{i})$ around the values $\tilde{w}(t_{i})$.
$PDw(t)\equiv{}_{\tilde{w}}\langle\psi,t|\psi,t\rangle_{\tilde{w}}\prod_{t_{i}=0}^{t}\frac{dw(t_{i})}{\sqrt{2\pi\lambda/dt}},$} (50)
[…]
The extension of collapse models into the context of relativistic quantum field theories is far from trivial, in part, because in such setting there is no fully satisfactory candidate for a position operator [68]  [^20] There are various other issues that arise such as the need to rely on relativistic versions of a smearing procedure ([60] for a discussion of those issues, and [10] for a proposal).. The problem is aggravated further when considering quantum field theory in curved spacetimes because particles are not fundamental objects (only quantum fields are [82]). Consequently, the notion of localization differs substantially from the particle-based picture: while particle localization is associated with position and momentum uncertainties, hence introducing naturally a length scale $r_{c}$, field localization is instead related to uncertainties in the field values and their conjugate momenta, suggesting a parameter with different physical characteristics (e.g. physical dimensions). Therefore, there is, at this time, no clear and straightforward connection between the two notions of state reduction required for these different contexts. Others, making use of different choices, obtain different results [54, 30]. The issue has been discussed in [15]. A similar issue arises for the collapse rate parameter $\lambda$, which is constrained by laboratory experiments but must be extrapolated to the early universe, a regime involving multiple phase transitions and poorly understood epochs. For these reasons, there is no compelling basis for a simple correspondence between CSL parameters defined at laboratory scales and those relevant in cosmological settings.
Nevertheless discussion above does not imply that a consistent phenomenological description in cosmology is unattainable. The cosmological setting is one in which, in practice, several of these conceptual difficulties can be bypassed, owing to the high degree of symmetry (even if only approximate) of the background spacetime and to the existence of preferred time parameters naturally associated with cosmic evolution. However, even then, multiplicity of options for the collapse operators $A$ exists (see, for instance, the discussion in [44] and the exploration of the results of some specific choices in [56, 14, 31]). The version of the CSL theory used in this manuscript is described by equations 49 and 50, using the field [^21] This proposal corresponds to taking the scalar field at each point as the collapse operator, and therefore replacing the sum over degrees of freedom $i$ by the spatial integral $\boldsymbol{\int\sqrt{h}d^{3}x}$. and its conjugate momentum as collapse operators $\hat{A}$. This particular formulation of the theory was introduced in [21] to account for the primordial power spectrum. That work considers various choices for the collapse operator $\hat{A}$ and found that the form of $\lambda$ that ensures an essencially scale invariant spectrum is required to become a function  [^22] In order to preserve scale invariance in the power spectrum, a specific functional form for the collapse rate $\lambda$ is chosen. of the field modes $k$. In particular, when the collapse operator is taken to be the rescaled field $a\delta\hat{\phi}$, the form that is needed is $\lambda(k)=\lambda_{0}k$. For typical values of the inflationary model parameters the value for the parameter $\lambda_{0}$ turns out to be $\lambda_{0}=10^{-19}\text{s}^{-1}$ a value which is interestingly close to the range of values contemplated for the parameters appearing in non-relativistic quantum mechanical versions of the theory [7]. On the other hand, in [43], a crucial modification of the parametrization $\lambda(k)=\lambda_{0}k$ was introduced (see (44); however, the value; however the value $\lambda_{0}=10^{-19}\text{s}^{-1}$ was retained) in order to address simultaneously the emergence of primordial structure and the problem of eternal inflation.
Appendix B Solving the CSL equations in the primordial power spectrum